OpenStudy (howard-wolowitz):

For the polynomial function .............. find all local and global extrema.

OpenStudy (howard-wolowitz):

\[f(x)=-2x^{2} -4x +16\]

OpenStudy (howard-wolowitz):

@AdamK

OpenStudy (howard-wolowitz):

A. No local extrema exist. B. The only extrema point is (−1, 18). C. The local and global extrema are: (0, -4), (0, 2) and (-1, 18). D. No global extrema exist.

OpenStudy (howard-wolowitz):

B. The only extrema point is (-1,18).

OpenStudy (howard-wolowitz):

@SolomonZelman

OpenStudy (howard-wolowitz):

@Michele_Laino

OpenStudy (howard-wolowitz):

Hi!

OpenStudy (michele_laino):

Hi!

OpenStudy (michele_laino):

please you have to compute the first derivative of f(x)

OpenStudy (howard-wolowitz):

Ok

OpenStudy (howard-wolowitz):

How?

OpenStudy (freckles):

If you don't know calculus, do you know what kinda graph you have? f(x)=ax^2+bx+c are well known to give what kinda graphs?

OpenStudy (michele_laino):

use this identity: \[\left( {{x^n}} \right)' = n{x^{n - 1}}\]

OpenStudy (howard-wolowitz):

you guys dont understand I have NO idea how to do this stuff

OpenStudy (freckles):

f(x)=ax^2+bx+c gives a graph of a parabola. The a tells us if the vertex will be a min or a max. If a>0, then the vertex will be a min. If a<0, then the vertex will be a max. Do you know how to find the vertex of a parabola ?

OpenStudy (howard-wolowitz):

yes hold on one sec

OpenStudy (howard-wolowitz):

ok so i did that then i replaced that for the x's and got the B one (-1,18)

OpenStudy (howard-wolowitz):

Michele what do you think about this?

OpenStudy (freckles):

Sounds great.

OpenStudy (michele_laino):

I think that your answer is correct!

OpenStudy (howard-wolowitz):

Awesome!

OpenStudy (howard-wolowitz):

I actually got it right. I did some math and did what u said and got it HURRRRAHHHHHH

OpenStudy (howard-wolowitz):

well thxs guys

OpenStudy (howard-wolowitz):

who gets the medal?

OpenStudy (michele_laino):

thanks!

OpenStudy (freckles):

Just so you know @Michele_Laino was trying to show you a calculus way. I don't know if you know calculus or not but it is also cute.

OpenStudy (michele_laino):

I think that @freckles deserves a medal

OpenStudy (howard-wolowitz):

I dont I'm only in Algebra 2

OpenStudy (freckles):

I don't care about medals. :p Go to calculus. I promise it will be fun.

OpenStudy (freckles):

I will give @Michele_Laino a medal. :)

OpenStudy (howard-wolowitz):

hey i got 19 more help freckles ok i give more medals

OpenStudy (howard-wolowitz):

For the following graph, determine if the graph represents a polynomial function. If it is a polynomial function, then determine the number of turning points and the least degree possible

OpenStudy (howard-wolowitz):

A. Yes, this graph represents a polynomial. There is one turning point and the least degree possible is two. B. Yes, this graph represents a polynomial. There are two turning points and the least degree possible is two. C. This is not a polynomial. There are no turning points. D. This is not a polynomial. The entire graph is on one side of the vertical axis.

OpenStudy (freckles):

Do you know what a turning point is?

OpenStudy (howard-wolowitz):

where the point on the graph rotates?

OpenStudy (freckles):

This is a point in where the graph changes direction.

OpenStudy (freckles):

like maybe it is going down and then it decides to go up (or vice versa) Where this happens, this is called a turning point

OpenStudy (howard-wolowitz):

so then it wouild be A?

OpenStudy (freckles):

Sounds great

OpenStudy (howard-wolowitz):

is that right?

OpenStudy (howard-wolowitz):

are you just agreeing with me or am i right?

OpenStudy (freckles):

Yep. It is a parabola. And parabola are polynomials. Parabolas have deg 2 and 1 turning point .

OpenStudy (howard-wolowitz):

dud AWESOMEEEEE

OpenStudy (howard-wolowitz):

Use the Remainder Theorem to find the remainder resulting from the division:

OpenStudy (howard-wolowitz):

\[(x ^{2}+5x-1) \div (x-1)\]

OpenStudy (freckles):

I'm going away from my computer now for a little bit after this next question so you might want to decide to post next one as a new question. But for this one... Just find when the bottom is zero...That is solve x-1=0 for x Then whatever you get for x there...plug into the x^2+5x-1 to find the remainder of (x^2+5x-1)/(x-1)

OpenStudy (howard-wolowitz):

A. 5 B. -7 C. -5 D. 0

OpenStudy (freckles):

Have you performed first step? That is solve x-1=0 for x. Then second step is: put that x into x^2+5x-1

OpenStudy (howard-wolowitz):

yes

OpenStudy (freckles):

cool and what did you get?

OpenStudy (howard-wolowitz):

I got -1

OpenStudy (freckles):

x-1=0 when x=?

OpenStudy (howard-wolowitz):

-2

OpenStudy (freckles):

x-1=0 plus 1 on both sides

OpenStudy (howard-wolowitz):

1 then

OpenStudy (freckles):

x-1=0 +1 +1 --- -- x+0=1 x=1 Yes the bottom is 0 when x is 1. So now take that 1 in plug into the top polynomial

OpenStudy (howard-wolowitz):

on both sides?

OpenStudy (howard-wolowitz):

i got 5

OpenStudy (freckles):

wht do you mean both sides?

OpenStudy (howard-wolowitz):

its A right

OpenStudy (freckles):

the top polynomial is x^2+5x-1 replace x with 1 and yes 5 is the remainder this works because: \[\frac{x^2+5x-1}{x-1}=Q(x)+\frac{R(x)}{x-1} \\ \text{ multiply } (x-1) \text{ on both sides } \\ x^2+5x-1=Q(x)(x-1)+R(x) \\ \text{pluggin \in x=1 will give us the remainder at x=1 } \\ 1^2+5(1)-1=Q(x)(1-1)+R(1) \\ 1+5-1=Q(x) \cdot 0 +R(1) \\ 5=R(1) \]

OpenStudy (howard-wolowitz):

Awesome well since u gotta go ill fan u and if u ever get back on ill contact u @frackles